40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 135     Entry time: Mon Jun 3 18:58:08 2013
Author: Koji 
Type: Optics 
Category: Configuration 
Subject: OMC final tests 

- QPD mount aligned, QPD output checked
  The spots are with 100um from the center of the diodes. [ELOG Entry (2nd photo)]

- TMS/FSR dependence on the PZT V
  Shows significant dependence on the PZT voltages
 
It seems that the curvartures get longer when the voltages are applied to the PZTs.
  The effect on these two PZTs are very similar. The dependence is something like
  (TMS/FSR) ~ 0.219 - 1e-5 V
  May cause resonance of the higher-order modes (like 13th order of the 45MHz sidebands) at a specific range of the PZTs.
  We can't change anything any more, but the impact needs to be assessed


- DC response of the PZTs [ELOG Entry]
  PZT voltages were swept. Observed multiple fringes during the sweep.
  The data to be analyzed.

- AC response of the PZTs [ELOG Entry]
  PZT1 and PZT2 well matched. The first resonance at 10kHz.

- Open loop TF of the servo
  The UGF more than ~30kHz.

- Cleaning of the main optics with First Contact
  Done. Visible scattering seen with an IR was reduced, but still exist.
  All four cavity mirrors have about the same level of scattering.
  Each scattering is a group of large or small bright spots.
  It's actually a bit difficult to resolve the bright spots with the IR viewer.

- Raw transmission: i.e. Ratio between the sum of the DCPD paths and the incident power
  May 8th (before the baking):      0.918
  May 8th (First Contact applied): 0.940 (improved)
  Jun 2nd (after the baking):         0.927 (worse)
  Jun 2nd (First Cotact applied):   0.964 (improved)

 

Date 2013/6/2 2013/6/2 2013/6/2
Condition  Before the cleaning  After the FC cleaning  After drag wiping
Input Power [mW]  39.8  38.4  38.4
REFLPD dark offset [V]  -0.0080  -0.0080  -0.0080
REFLPD locked [V]  0.048  0.0437  0.046
REFLPD unlocked [V]  6.41  6.39 6.37
       
 Transmitted Power to DCPD1 (T) [mW]  18.8  18.8  18.8
 Transmitted Power to DCPD2 (R) [mW]  18.1  18.2  18.2
 FM2 transmission [mW]  -  -  -
 CM1 transmission [mW]  0.200  0.193  0.198
 CM2 transmission [mW]  0.204  0.204  0.205
 Input BS transmission [mW]  0.260  0.228  0.245
       
 Cavity Finesse 396.9  403.79  403.79
       
 Junk Light Power (Pjunk) [mW]  0.303  0.302  0.317
 Coupled beam power (Pcouple) [mW]  39.50  38.10  38.08
 Mode Matching (Pcouple/Pin) [mW]  0.992  0.992  0.992
 Cavity reflectivity in power  0.00112  0.000211  0.000206
 Loss per mirror [ppm]  111  35.9  34.8
 Cavity transmission for TEM00 carrier
 0.934  0.971  0.972

 

- TMS/FSR/Finesse change before/after cleaning [ELOG Entry]
  Just a small change from the parameters before the bake.
  No quantitative difference.

  Method:
  BB EOM produces the AM sidebands together with the PM sidebands.
  Ideally, the PM sidebands does not produce the signal at the transmission, the output is dominated by the AM component.
  This is only true when there is no lock offset. In reality the curve is contaminated by the PM-AM conversion by the
  static offset or dynamic deviation of the locking point. So I had to take the central part of the TF and check the
  dependence of the fit region and the finesse.

  Before the cleaning: Finesse 396.9
  After the cleaning: Finesse 403.8


To Do

- Placement of the DCPD housings
- Through-put test with DCPDs
- Transmission dependence on the incident power
  (although the max incident is limited to ~35mW)

- Application of the first contact for the surface protection

ELOG V3.1.3-